Grade 12thMechanics

Hey man throw about maximum horizontal distance of 80 metre calculate the maximum height reached

Sayyad mushrraf
8 Years agoGrade 12th

1 Answer

Arun
8 Years ago
Dear sayyad
 
We know that R = u2 * Sin 2θ / g.
As, the range is maximum, so, θ = 450
So, 80 = u2 /g.
Now, H = height reached by ball =  u2 Sin2θ / 2g.
Putting value of θ = 45, u2 / g = 80, we get
H = 80/4  = 20m
 
Regards
Arun (askIITians forum expert)